A charm quark has a charge of approximately 1. 5.33 x 10-20 C 2. 1.07 x 10-19 3. 1.60 10-19 4. 2.40 x 10-19 C
Answer #1

Charge on Charm quark is +(2/3)e. Where e is elementary charge unit (magnitude of charge on electron or proton).

e= 1.6 × 10-19 C

So Charge on Charm quark is (2/3)×1.6×10-19

= 1.067 × 10-19

= 1.07 × 10-19

correct answer 2.